Again, please use the same data as in the previous question. It will be repeated for your convenience each time.

The following is production data:

Year Waffles
(Quantity)
Waffles
(Price)
Orange Juice
(Quantity)
Orange Juice
(Price)
Toaster Ovens
(Quantity)
Toaster Ovens
(Price)
2010 1000 $1.20 400 $5 280 $22
2011 1200 $1.30 410 $6 300 $25
2012 1500 $1.40 420 $7 300 $28

The above, like I said, is production data. Consumption data is as follows: the average citizen consumes 80 waffles, 20 units of orange juice and 10 toaster ovens. 2010 is the base year for all calculations.

Carefully following all directions for entering your answer, tell me the 2012 CPI value for this economy.
Answer #1

Answer : CPI = (Cost basket at current price level / Cost basket at base year price level) * 100

In 2012 :

CPI = [{($1.40 * 80) + ($7 * 20) + ($28 * 10)} / {($1.20 * 80) + ($5 * 20) + ($22 * 10)}] * 100

=> CPI = [(112 + 140 + 280) / (96 + 100 + 220) ] * 100

=> CPI = (532 / 416) * 100 = 1.2788 * 100

=> CPI =127.88

Therefore, here the CPI of 2012 is 127.88 .